OpenStudy (babynini):

Use squeeze theorem to show that lim sqroot (x^3+x^2)sin(pi/x)=0 x-> 0 illustrate by graphing the functions f,g,h on the same screen.

OpenStudy (babynini):



OpenStudy (babynini):

@zepdrix

OpenStudy (babynini):

@Astrophysics

zepdrix (zepdrix):

Which piece is causing trouble? \(\large\rm \sqrt{x^3+x^2}\) or the \(\large\rm \sin\left(\frac{\pi}{x}\right)\)? :)

OpenStudy (babynini):

I'm not sure where to start xD

zepdrix (zepdrix):

OpenStudy lag :O( lameeeee

OpenStudy (babynini):

hmm?

OpenStudy (anonymous):

As we know, sin of any number will always be between -1 and 1, inclusive.

OpenStudy (anonymous):

\[-1\le \sin \frac{ \pi }{ x }\le1\]

OpenStudy (anonymous):

now we multiply whole thing by \[\sqrt{x ^{3}+x ^{2}}\]

OpenStudy (babynini):

multiply the whole sin(pi/x)?

zepdrix (zepdrix):

Plug in x=0, which part of the function is causing a problem?

OpenStudy (anonymous):

yes like this

OpenStudy (anonymous):

its a squezz theorm @zepdrix

OpenStudy (anonymous):

ok i ll do it step by step u ll understand let me finish

OpenStudy (babynini):

(@zepdrix the sin part?) okay @ayeshaafzal221 :)

zepdrix (zepdrix):

..

OpenStudy (anonymous):

\[-(\sqrt{x ^{3}+x ^{2}})\le \sqrt{x ^{3}+x ^{2}} \times \sin \frac{ \pi }{ x }\le \sqrt{x ^{3}+x ^{2}}\]

zepdrix (zepdrix):

Yes, so start by putting bounds on the bad part. Sine function is bound by -1 and 1, yes? :)

OpenStudy (babynini):

Yeah

OpenStudy (anonymous):

With an inequality like this, you can add "lim x->0" to all 3 parts and the expression will still be valid.

zepdrix (zepdrix):

\[\large\rm -1\le \sin\left(\frac{\pi}{x}\right)\le 1\]

OpenStudy (babynini):

Okay, @zepdrix next what? :)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}-(\sqrt{x ^{3}+x ^{2}})\le \lim_{x \rightarrow 0} \sqrt{x ^{3}+x ^{2}} \times \sin \frac{ \pi }{ x }\le \lim_{x \rightarrow 0} \sqrt{x ^{3}+x ^{2}}\]

OpenStudy (anonymous):

now you solve left hand side and then right hand side

zepdrix (zepdrix):

:p

OpenStudy (anonymous):

so left hand side is \[\lim_{x \rightarrow 0} -(\sqrt{0^{3} +0^{2}} =-\sqrt{0}=0\]

OpenStudy (anonymous):

i am sub x=0

OpenStudy (anonymous):

now right hand side \[\lim_{x \rightarrow 0}\sqrt{0^{3}+0^{2}}=0\]

OpenStudy (anonymous):

thus the limit of the original expression is less than and equal to 0, and greater and equal to 0, then its limit must be 0.

OpenStudy (babynini):

ooo so, to answer the question, to I have to write out that proof?

OpenStudy (anonymous):

if u didnt understood any step let me know and i know what you confuse about whenever there is squeez theorm there are restriction that u must know before hand for cos , tan and sin , u can google em

OpenStudy (babynini):

The question says to "use notation of the squeeze theorem"

OpenStudy (anonymous):

yes :) thats what squeex theorm is

OpenStudy (babynini):

Okay, because I don't remember multiplying things in doing limits in class o.o

OpenStudy (anonymous):

ur question clearly says use squeez theorm to show that expression =0

OpenStudy (babynini):

Yeah. Usually we approach the limit from the left and right and see if they match up o.0 or calculate it as it approaches 0

OpenStudy (anonymous):

yup thats what i did

OpenStudy (anonymous):

see these domain and range u must know by heart its really useful

OpenStudy (anonymous):

in ur case just remember cos , sin and tan

OpenStudy (babynini):

ahh ok

OpenStudy (babynini):

I am beginning to understand.

OpenStudy (anonymous):

good to know just tag if u have any other problem :)

OpenStudy (babynini):

ok thanks!!

OpenStudy (babynini):

i) -1 </= sin (pi/x) </= 1 ii) \[-\sqrt{x^3+x^2}\le \sqrt{x^3+x^2}\sin(\frac{ \pi }{ x })\le \sqrt{x^3+x^2}\] iii) we know that \[\lim \sqrt{x^3+^2} = 0 \space and \space \lim(-\sqrt{x^3+x^2}=0\] Taking f(x)= -sq( x^3+x^2), g(x)= sq(x^3+x^2)sin(pi/x) and, h(x) = sq(x^3+x^2), in squeeze theorem, we conclude: lim sq(x^3+x^2)sin(pi/)=0 (as approaches 0)

OpenStudy (babynini):

*sin(pi/x)

OpenStudy (babynini):

@ayeshaafzal221 does that all look correct?

OpenStudy (anonymous):

yes @Babynini well done u got a hang of it :)

OpenStudy (babynini):

Thank you! :)